# Magnetic field amplification during a turbulent collapse

### v2

@@ 3.3 Critical vorticity
Figure 5: Evolution of $k_{\nabla\cdot\bm{u}}/k_{0}$ (upper panel) and $\mathcal{B}_{\rm rms}$ (lower panel) for $\zeta=0.1$ (red), $0.5$ (orange), $0.9$ (green), $0.95$ (blue), and $1$ (black). Runs 15–18 and Run 23.
@@
Figure 6: $k_{\omega}/k_{0}$ (upper panel) and $\mathcal{B}_{\rm rms}$ (lower panel) for $1$ (dotted black), $0.99$ (solid black), $0.98$ (blue), $0.97$ (green), $0.96$ (orange), and $\zeta=0.95$ (red). Runs 18–23.
@@ 3.4 Effect of scale separation
Figure 8: $k_{\omega}/k_{0}$ for hydrodynamic runs with $\zeta=1$, $\mbox{\rm Re}_{\rm M}=900$, and different values of $k_{0}$. For $k_{0}/k_{1}=10$, we also compare with the magnetic run with $\mbox{\rm Re}_{\rm M}=900$. Runs 24–27.
@@ 3.5 Growth of vorticity
For an isothermal gas, there is no baroclinic term, which would be the main agent for producing vorticity in nonisothermal flows. There is also no rotation nor shear, both of which could lead to vorticity generation (Del Sordo & Brandenburg, 2011; Elias-López et al., 2023, 2024). There remain only three possibilities for driving or amplifying vorticity: (i) through viscosity via gradients of the velocity divergence being inclined against density gradients, (ii) through magnetic driving or magnetically assisted vorticity production (Brandenburg & Scannapieco, 2025), and (iii) through nonlinearity.
@@ 3.6 Spectral evolution
As already demonstrated in Brandenburg & Ntormousi (2022), the collapse dynamics does not affect the magnetic energy spectra significantly. At length scales above the Jeans length, the collapse does lead to a growth of the compressive part of the kinetic energy spectra and even a growth of magnetic energy, but this is associated with the compression itself and is not a consequence of a dynamo; see Figure 9(b) of Brandenburg & Ntormousi (2022).
@@ 3.7 Instantaneous growth rate
$-\langle\bm{J}\cdot(\bm{u}\times\bm{B})\rangle=\langle J_{i}u_{j}(A_{i,j}-A_{j,i}\rangle\equiv W_{\rm L}^{\rm 2D}+W_{\rm L}^{\rm 3D}.$ (16)
@@ 3.8 Visualizations
In Figure 15, we present visualizations of $\mathcal{B}_{z}$, $\omega_{z}/u_{0}k_{0}$, and ${\bm{\nabla}}\cdot\bm{u}/u_{0}k_{0}$ for Run 37 at early and late times. There is no significance in us having chosen the $z$ component of $\bm{B}$ and $\bm{\omega}$; all three components are statistically equivalent.
@@ 4 Comparison with previous work
In our earlier paper (Brandenburg & Ntormousi, 2022), we simulated gravitational collapse using numerical simulations of decaying turbulence in a Jeans-unstable domain at a resolution of $2048^{2}$ mesh points. We only found a weak increase of the magnetic field with time. Given the knowledge of the collapse time from the simulations, i.e., the time when the singularity was reached, we can replace the pressure-less free-fall time by the actual collapse time and express the evolution of the rms magnetic field in comoving coordinates. This allows us to see whether the growth in the old simulations is close to exponential in comoving coordinates during any time interval.
[…]
@@ 5 Conclusions
In this work we approached the problem of dynamo action during gravitational collapse by employing supercomoving coordinates. This is a significant change of paradigm with respect to previous simulation work (e.g., Sur et al., 2010, 2012; Federrath et al., 2011b; Brandenburg & Ntormousi, 2022) which was limited by the shrinking dynamical range during the collapse; see Section 4. In supercomoving coordinates we can look for exponential growth of the magnetic field, which is a clear signature of a dynamo. This allows us to surpass the other obstacle faced by previous work, which is characterizing dynamos in unsteady flows.
@@
Irshad P et al. (2025) considered forced turbulence, as opposed to our study of decaying turbulence. Therefore, in their models, the magnetic field could always be sustained, but the source of the driving remains unclear. The superexponential growth that they reported, however, it still recovered in our decay simulations, unless the free-fall time is longer than the turnover time of the turbulence. In that case, the growth is actually subexponential, but this is primarily a consequence of the decay of the turbulence.
@@
Earlier work on turbulent collapse and dynamo action has suggested that the collapse drives turbulence and enhanced it (Sur et al., 2012; Xu & Lazarian, 2020; Hennebelle, 2021). Our work casts doubt on this interpretation, because of two aspects. First, the collapse dynamics reduces the effective nonlinearity, resulting in stronger apparent field amplification by the turbulence, and second, there can be generation of vorticity both from viscosity and from the magnetic field itself, but this is not directly related to the collapse. It should therefore be checked, whether these two factors could have contributed to the earlier findings of collapse-driven turbulence. In this context, the fact that we do not solve the Poisson equation for self-gravity but treat the collapse as a homogeneous flow through the change of coordinates could be a difference worth investigating.
As explained in Section 4, the transformation to supercomoving coordinates may also help analyzing existing simulations in physical coordinates. We argue that for homogeneous collapse simulations that do not utilize supercomoving coordinates, it is still useful to express such results in terms of comoving quantities and conformal time, because they might display exponential magnetic field growth that would be the perhaps strongest indication of dynamo action so far.
Our work has applications not just to interstellar clouds and primordial star formation (e.g., Schleicher et al., 2009; Hirano & Machida, 2022; Sharda et al., 2020), but also to larger cosmological scales. Our results show that small amounts of vorticity might suffice to produce dynamo action even in decaying turbulence which, we argue, is also relevant to gravitational collapse. This consideration is important for understanding magnetism in protohalos before the first stars form and their feedback drives sufficient turbulence for dynamo action (e.g., Schleicher et al., 2010).
@@
We thank Fabio Del Sordo for helpful discussions on vorticity generation, Pallavi Bhat for suggesting an explanation for the subexponential growth for $\mathcal{S}\ll 1$, and the referee for a detailed assessment of our work. We also acknowledge inspiring discussions with the participants of the program on “Turbulence in Astrophysical Environments” at the Kavli Institute for Theoretical Physics in Santa Barbara. This research was supported in part by the Swedish Research Council (Vetenskapsrådet) under Grant No. 2019-04234, the National Science Foundation under Grants No. NSF PHY-2309135, AST-2307698, and NASA Awards 80NSSC22K0825 and 80NSSC22K1265. E.N. acknowledges funding from the Italian Ministry for Universities and Research (MUR) through the “Young Researchers” funding call (Project MSCA 000074). We acknowledge the allocation of computing resources provided by the Swedish National Allocations Committee at the Center for Parallel Computers at the Royal Institute of Technology in Stockholm. Software and Data Availability. The source code used for the simulations of this study, the Pencil Code (Pencil Code Collaboration et al., 2021), is freely available on https://github.com/pencil-code. The simulation setups and corresponding input and reduced output data are freely available on http://doi.org/10.5281/zenodo.15693287 (catalog http://doi.org/10.5281/zenodo.15693287).

### v3

@@ 3.3 Critical vorticity
Figure 5: Evolution of $k_{\nabla\cdot\bm{u}}/k_{0}$ (upper panel) and $\mathcal{B}_{\rm rms}$ (lower panel) for $\zeta=0.1$ (red), $0.5$ (orange), $0.9$ (green), $0.95$ (blue), and 11 (black). Runs 15–18 and Run 23.
@@
Figure 6: $k_{\omega}/k_{0}$ (upper panel) and $\mathcal{B}_{\rm rms}$ (lower panel) for 11 (dotted black), $0.99$ (solid black), $0.98$ (blue), $0.97$ (green), $0.96$ (orange), and $\zeta=0.95$ (red). Runs 18–23.
@@ 3.4 Effect of scale separation
Figure 8: $k_{\omega}/k_{0}$ for hydrodynamic runs with $\zeta=1$, $\mbox{\rm Re}_{\rm M}=900$, and different values of $k_{0}$. Runs 24–27.
@@ 3.5 Growth of vorticity
For an isothermal gas, there is no baroclinic term, which would be the main agent for producing vorticity in nonisothermal flows. There is also no rotation or shear, both of which could lead to vorticity generation (Del Sordo & Brandenburg, 2011; Elias-López et al., 2023, 2024). There remain only three possibilities for driving or amplifying vorticity: (i) through viscosity via gradients of the velocity divergence being inclined against density gradients, (ii) through magnetic driving or magnetically assisted vorticity production (Brandenburg & Scannapieco, 2025), and (iii) through nonlinearity.
@@ 3.6 Spectral evolution
As already demonstrated in Brandenburg & Ntormousi (2022), the collapse dynamics do not affect the magnetic energy spectra significantly. At length scales above the Jeans length, the collapse does lead to a growth of the compressive part of the kinetic energy spectra and even a growth of magnetic energy, but this is associated with the compression itself and is not a consequence of a dynamo; see Figure 9(b) of Brandenburg & Ntormousi (2022).
@@ 3.7 Instantaneous growth rate
$\langle\bm{J}\cdot(\bm{u}\times\bm{B})\rangle=\langle J_{i}u_{j}(A_{i,j}-A_{j,i}\rangle\equiv W_{\rm L}^{\rm 2D}+W_{\rm L}^{\rm 3D}.$ (16)
@@ 3.8 Visualizations
In Figure 15, we present visualizations of $\mathcal{B}_{z}$, $\omega_{z}/u_{0}k_{0}$, and ${\bm{\nabla}}\cdot\bm{u}/u_{0}k_{0}$ for Run 37 at early and late times. There is no significance in our having chosen the $z$ component of $\bm{B}$ and $\bm{\omega}$; all three components are statistically equivalent.
@@ 4 Comparison with previous work
In our earlier paper (Brandenburg & Ntormousi, 2022), we simulated gravitational collapse using numerical simulations of decaying turbulence in a Jeans-unstable domain at a resolution of $2048^{2}$ mesh points. We only found a weak increase in the magnetic field with time. Given the knowledge of the collapse time from the simulations, i.e., the time when the singularity was reached, we can replace the pressureless freefall time by the actual collapse time and express the evolution of the rms magnetic field in comoving coordinates. This allows us to see whether the growth in the old simulations is close to exponential in comoving coordinates during any time interval.
The result is shown in Figure 16, where we computed the conformal time and scale factor numerically based on Equation (1). Here we used the empirical value of $t_{\rm ff}\approx 2.016/c_{\rm s}k_{1}$, which yields $s\approx 0.78\,c_{\rm s}k_{0}$, and thus, since $u_{0}/c_{\rm s}=0.19$ and $k_{0}/k_{1}=10$, we have $\mathcal{S}\approx 0.4$; see Table LABEL:Tsummary, where it is called Run 39. The physical values of the magnetic field computed by Brandenburg & Ntormousi (2022) are denoted by $\mathcal{B}_{\mathrm{ph}}$. We also plot the comoving values $a^{2}\mathcal{B}_{\mathrm{ph}}$ both versus physical and conformal time. Here, the $a(t)$ and the conformal time have been computed from Equations (1) and (2). Although there is a steady increase of $\mathcal{B}_{\rm rms}$, Figure 16(b) shows that the comoving magnetic field does not follow an exponential growth in conformal time, except for a very early time interval $0<tu_{0}k_{0}\mathrel{\mathchoice{\vbox{\offinterlineskip\halign{\hfil$\displaystyle#$\hfil\cr<\cr\sim\cr}}}{\vbox{\offinterlineskip\halign{\hfil$\textstyle#$\hfil\cr<\cr\sim\cr}}}{\vbox{\offinterlineskip\halign{\hfil$\scriptstyle#$\hfil\cr<\cr\sim\cr}}}{\vbox{\offinterlineskip\halign{\hfil$\scriptscriptstyle#$\hfil\cr<\cr\sim\cr}}}}0.4$.
@@ 5 Conclusions
In this work we approached the problem of dynamo action during gravitational collapse by employing supercomoving coordinates. This is a significant change of paradigm with respect to previous simulation work (e.g., Sur et al., 2010, 2012; Federrath et al., 2011b; Brandenburg & Ntormousi, 2022), which was limited by the shrinking dynamical range during the collapse; see Section 4. In supercomoving coordinates, we can look for exponential growth of the magnetic field, which is a clear signature of a dynamo. This allows us to surpass the other obstacle faced by previous work, which is characterizing dynamos in unsteady flows.
@@
Irshad P et al. (2025) considered forced turbulence, as opposed to our study of decaying turbulence. Therefore, in their models, the magnetic field could always be sustained, but the source of the driving remains unclear. The superexponential growth that they reported, however, was still recovered in our decay simulations, unless the freefall time is longer than the turnover time of the turbulence. In that case, the growth is actually subexponential, but this is primarily a consequence of the decay of the turbulence.
@@
Earlier work on turbulent collapse and dynamo action has suggested that the collapse drives turbulence and enhances it (Sur et al., 2012; Xu & Lazarian, 2020; Hennebelle, 2021). Our work casts doubt on this interpretation due to two aspects. First, the collapse dynamics reduces the effective nonlinearity, resulting in stronger apparent field amplification by the turbulence, and second, there can be generation of vorticity both from viscosity and from the magnetic field itself, but this is not directly related to the collapse. It should therefore be checked whether these two factors could have contributed to the earlier findings of collapse-driven turbulence. In this context, the fact that we do not solve the Poisson equation for self-gravity but treat the collapse as a homogeneous flow through the change of coordinates could be a difference worth investigating.
As explained in Section 4, the transformation to supercomoving coordinates may also help analyze existing simulations in physical coordinates. We argue that for homogeneous collapse simulations that do not utilize supercomoving coordinates, it is still useful to express such results in terms of comoving quantities and conformal time, because they might display exponential magnetic field growth that would be perhaps the strongest indication of dynamo action so far.
Our work has applications not just to interstellar clouds and primordial star formation (e.g., Schleicher et al., 2009; Hirano & Machida, 2022; Sharda et al., 2020), but also to larger cosmological scales. Our results show that small amounts of vorticity might suffice to produce dynamo action even in decaying turbulence, which, we argue, is also relevant to gravitational collapse. This consideration is important for understanding magnetism in protohalos before the first stars form and their feedback drives sufficient turbulence for dynamo action (e.g., Schleicher et al., 2010).
@@
We thank Fabio Del Sordo for helpful discussions on vorticity generation, Pallavi Bhat for suggesting an explanation for the subexponential growth for $\mathcal{S}\ll 1$, and the referee for a detailed assessment of our work. We also acknowledge inspiring discussions with the participants of the program on “Turbulence in Astrophysical Environments” at the Kavli Institute for Theoretical Physics in Santa Barbara. This research was supported in part by the Swedish Research Council (Vetenskapsrådet) under grant No. 2019-04234, the National Science Foundation under grants No. NSF PHY-2309135, AST-2307698, AST-2408411, and NASA Award 80NSSC22K0825. E.N. acknowledges funding from the Italian Ministry for Universities and Research (MUR) through the “Young Researchers” funding call (Project MSCA 000074). We acknowledge the allocation of computing resources provided by the Swedish National Allocations Committee at the Center for Parallel Computers at the Royal Institute of Technology in Stockholm. Software and Data Availability. The source code used for the simulations of this study, the Pencil Code (Pencil Code Collaboration et al., 2021), is freely available on https://github.com/pencil-code. The simulation setups and corresponding input and reduced output data are freely available on http://doi.org/10.5281/zenodo.15693287 (catalog http://doi.org/10.5281/zenodo.15693287).